If Leah is 6 years older than her sister, Sue, and John is 5 years older than Leah, and the total of their ages is 41. Then how old is Sue? This question is answered only john in the equation was forgotten. Please answer showing an equation to figure this out.
Answer by rapaljer(4671)

Let L = Leah's age
S= Sue's age
J= John's age

L = S+6
J= L + 5 = (S+6) + 5 = S + 11

L + S + J = 41 Substitute everything to obtain an equation in S alone:
(S+6) + S + (S+ 11) = 41
3S + 17 = 41

Subtract 17 from each side:
3S + 17 = 41 - 17
3S = 24
S= 8 Sue
L= 8+6 = 14 Leah
J= 6+11 = 17 John

Check: 8 + 14 + 17 = 41

R^2 at SCC